$SPAD/src/input romanpolynomials.input

Timothy Daly December 3, 2016

Abstract

Contents

2

— * —

)set break resume
)sys rm -f romanpolynomials.output
)spool romanpolynomials.output
)set message test on
)set message auto off
)clear all
--S 1 of 14
(x-4::ROMAN)
--R
--R
--R   (1)  VI
--R                                               Type: Polynomial(RomanNumeral)
--E 1
--S 2 of 14
x
--R
--R
--R   (2)  x
--R                                                            Type: Variable(x)
--E 2
--S 3 of 14
x::ROMAN
--R
--R
--R   (3)  X
--R                                                           Type: RomanNumeral
--E 3
--S 4 of 14
a::ROMAN
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   Improper character in Roman numeral:
--R   #\A
--R
--R   Continuing to read the file...
--R
--E 4
--S 5 of 14
a::ROMAN
--R

--R
--RDaly Bug
--R   >> Error detected within library code:
--R   Improper character in Roman numeral:
--R   #\A
--R
--R   Continuing to read the file...
--R
--E 5
--S 6 of 14
d::ROMAN
--R
--R
--R   (4)  D
--R                                                           Type: RomanNumeral
--E 6
--S 7 of 14
M::ROMAN
--R
--R
--R   (5)  M
--R                                                           Type: RomanNumeral
--E 7
--S 8 of 14
(y-4::ROMAN)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   Improper character in Roman numeral:
--R   #\Y
--R
--R   Continuing to read the file...
--R
--E 8
--S 9 of 14
x:POLY(ROMAN):='x
--R
--R
--R   (6)  X
--R                                               Type: Polynomial(RomanNumeral)
--E 9
--S 10 of 14
(x-4::ROMAN)
--R

--R
--R   (7)  VI
--R                                               Type: Polynomial(RomanNumeral)
--E 10
--S 11 of 14
(x-4)
--R
--R
--R   (8)  VI
--R                                               Type: Polynomial(RomanNumeral)
--E 11
--S 12 of 14
x::Symbol::Polynomial(ROMAN)
--R
--R
--RDaly Bug
--R   Cannot convert from type Polynomial(RomanNumeral) to Symbol for
--R      value
--R   X
--R
--E 12
--S 13 of 14
x:=monomial(1,monomial(1,'x)$IndexedExponents(Symbol))$Polynomial(ROMAN)
--R
--R
--R   (9)  x
--R                                               Type: Polynomial(RomanNumeral)
--E 13
--S 14 of 14
(x-5::ROMAN)^2
--R
--R
--R          2
--R   (10)  x  - X x + XXV
--R                                               Type: Polynomial(RomanNumeral)
--E 14
)spool
)lisp (bye)

———-

References^{[1] nothing}

6